# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ Appendix A Algorithmic details
where $N=nm-mp-nq-(n+m)d-1$ is the effective degrees of freedom of the model, that is the difference between the number of observations and the number of parameters to be estimated. This can be computed a posteriori or iteratively refined during the optimization substituting $\hat{\boldsymbol{\mu}}$ with $\boldsymbol{\mu}^{t}$.
@@
$D_{\alpha}(y,\mu)=2w\bigg{[}y\log\frac{y}{\mu}-(y+\alpha)\log\frac{y+\alpha}{\mu+\alpha}\bigg{]},\quad V_{\alpha}(\mu)=w\mu(1+\mu/\alpha).$
Since the inflation parameter $\alpha>0$ is typically unknown, we need to estimate it from the data. A common choice in the literature is to consider the moment estimator
$\hat{\alpha}=\frac{\displaystyle\bigg{[}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{ij}\hat{\mu}_{ij}^{2}\bigg{]}}{\displaystyle\bigg{[}\sum_{i=1}^{n}\sum_{j=1}^{m}w_{ij}\big{\{}(y_{ij}-\hat{\mu}_{ij})^{2}-\hat{\mu}_{ij}\big{\}}\bigg{]}}=\frac{\mathbf{1}_{n}^{\top}(\mathbf{W}*\hat{\boldsymbol{\mu}}*\hat{\boldsymbol{\mu}})\mathbf{1}_{m}}{\mathbf{1}_{n}^{\top}\big{[}\mathbf{W}*\{(\mathbf{Y}-\hat{\boldsymbol{\mu}})^{2}-\hat{\boldsymbol{\mu}}\}\big{]}\mathbf{1}_{m}}.$
@@
Initialization details
As explained in Section 3.4 of the main paper, in OLS–SVD initialization, we transform the data according to the perturbed link transformation $y_{\varepsilon,ij}=g_{\varepsilon}(y_{ij})$, which depends on the control parameter $\varepsilon>0$, which is typically a small constant. In particular, we consider the following transformation
$g_{\varepsilon}(y)=\begin{cases}g(\varepsilon)\,\mathbb{I}_{\{y=0\}}+g(1-\varepsilon)\,\mathbb{I}_{\{y=1\}}&\text{for binary families},\\ g(\varepsilon)\,\mathbb{I}_{\{y=0\}}+g(y)\,\mathbb{I}_{\{y>0\}}&\text{for count families},\\ g(y)&\text{otherwise},\end{cases}$
where $\mathbb{I}_{A}$ is the indicator function of set $A$, being equal to $1$ when $A$ is satisfied and 00 otherwise.
Appendix B Simulation setting details
@@
  • NMF+: we use the NNMF() function in the NNLM package (Lin and Boutros,, 2020), and we specify the following options: k = $d$, alpha = 1, beta = 1, n.threads = 4, method = "lee", loss = "mkl", max.iter = 2000.
@@
  • NBWaVE: we use the newFit() function in the NewWave package (Agostinis et al.,, 2022), and we specify the following options: K = $d$, commondispersion = TRUE, maxiter_optimize = 500, stop_epsilon = 1e-04, children = 4.
@@
  • SGD: we use the cpp.fit.bsgd() function in the sgdGMF package, and we specify the following options: ncomp = $d$, familyname = "poisson", linkname = "log", lambda = c(0,0,1,0), maxiter = 500, rate0 = 0.01, size1 = 100, size2 = 20, eps = 1e-08, nafill = 1, tol = 1e-05, damping = 1e-03.
It is worth mentioning that the original R implementation of the AIRWLS and Newton methods can be found in the gmf package by Kidziński et al., (2022). However, such an implementation does not permit the inclusion of gene-specific intercepts and covariate effects, say $\boldsymbol{\gamma}_{i}^{\top}\boldsymbol{z}_{j}$ in our notation, and, also, it does not allow for parallel computing in Windows operating systems. Therefore, we performed the benchmarking experiments using the R/C++ implementation in the proposed sgdGMF package.
@@
Appendix C Supplementary figures
Supplementary Fig. 6: Comparison of SGD using block-subsampling with NBWaVE (Agostinis et al.,, 2022) and glmPCA (Townes et al.,, 2019) on the Arigoni dataset (Arigoni et al.,, 2024). All methods use a matrix rank of 15, which was suggested by model selection criteria. A) tSNE embeddings show no clear differences between all methods. B) SGD is orders of magnitude faster compared to the other methods, and has a lower out-of-sample deviance prediction error when predicting missing values. This is probably due to NBWaVE and glmpca requiring imputation of missing values prior to computation of the latent structure, while SGD can deal with missing values internally. C) The mean cell-line cluster purity is also similar for all methods. SGD gives similar results on all levels when using the Poisson and negative binomial family on the Arigoni dataset. Therefore, SGD is a fast alternative to NBWaVE and glmPCA , while having similar performance.
Supplementary Fig. 7: Eigenvalues of the latent directions ordered in descending values for the TENxBrainData (Lun and Morgan,, 2023). This plot, also called an elbow-plot can be used for model selection, by looking for an elbow in the curve. Here, 10 latent factors were chosen to be used in the case study.

### v2

@@ Appendix B Additional algorithmic details
where N=nmmpnq(n+m)d1N=nm-mp-nq-(n+m)d-1 is the effective degrees of freedom of the model, that is the difference between the number of observations and the number of parameters to be estimated. This can be computed a posteriori or iteratively refined during the optimization substituting $\hat{\boldsymbol{\mu}}$ with $\boldsymbol{\mu}^{t}$.
@@
$D_{\alpha}(y,\mu)=2w\bigg[y\log\frac{y}{\mu}-(y+\alpha)\log\frac{y+\alpha}{\mu+\alpha}\bigg],\quad\nu_{\alpha}(\mu)=w\mu(1+\mu/\alpha),$
where $\alpha$ is the shape parameter of the Negative Binomial family. Since the shape parameter $\alpha>0$ is typically unknown, we need to estimate it from the data. A common choice in the literature is to consider the moment estimator
$\hat{\alpha}=\frac{\displaystyle\bigg[\sum_{i=1}^{n}\sum_{j=1}^{m}w_{ij}\hat{\mu}_{ij}^{2}\bigg]}{\displaystyle\bigg[\sum_{i=1}^{n}\sum_{j=1}^{m}w_{ij}\big\{(y_{ij}-\hat{\mu}_{ij})^{2}-\hat{\mu}_{ij}\big\}\bigg]}=\frac{\mathbf{1}_{n}^{\top}(\mathbf{W}*\hat{\boldsymbol{\mu}}*\hat{\boldsymbol{\mu}})\mathbf{1}_{m}}{\mathbf{1}_{n}^{\top}\big[\mathbf{W}*\{(\mathbf{Y}-\hat{\boldsymbol{\mu}})^{2}-\hat{\boldsymbol{\mu}}\}\big]\mathbf{1}_{m}},$
@@
Appendix C Simulation setting details
@@
NMF+: we use the NNMF() function in the NNLM package (Lin and Boutros,, 2020), and we specify the following options: k = $d$, alpha = 1, beta = 1, n.threads = 4, method = "lee", loss = "mkl", max.iter = 1000.
@@
NBWaVE: we use the newFit() function in the NewWave package (Agostinis et al.,, 2022), and we specify the following options: K = $d$, commondispersion = TRUE, maxiter_optimize = 200, stop_epsilon = 1e-05, children = 4.
GFM-AM: we use the gfm() function in the GFM package (Liu et al.,, 2023), and we specify the following options: types = "poisson", q = $d$, offset = FALSE, dc_eps = 1e-05, maxIter = 200, algorithm = "AM".
GFM-VEM: we use the gfm() function in the GFM package (Liu et al.,, 2023), and we specify the following options: types = "poisson", q = $d$, offset = FALSE, dc_eps = 1e-05, maxIter = 200, algorithm = "VEM".
COAP: we use the COAP_RR() function in the COAP package (Liu and Zhong,, 2024), and we specify the following options: Z = $\mathbf{X}$, q = $d$, epsELBO = 1e-05, maxIter = 100, joint_opt_beta = FALSE, fast_svd = FALSE.
@@
aSGD: we use the cpp.fit.bsgd() function in the sgdGMF package, and we specify the following options: ncomp = $d$, familyname = "poisson", linkname = "log", lambda = c(0,0,1,0), maxiter = 500, rate0 = 0.01, size1 = 100, size2 = 20, eps = 1e-08, nafill = 1, tol = 1e-05, damping = 1e-03.
It is worth mentioning that the original R implementation of the AIRWLS and Newton methods can be found in the gmf package (Kidziński et al.,, 2022). However, such an implementation does not permit the inclusion of gene-specific intercepts and covariate effects, say $\boldsymbol{\gamma}_{i}^{\top}\boldsymbol{z}_{j}$ in our notation, and also it does not allow for parallel computing in Windows operating systems. Therefore, we performed the benchmarking experiments using the R/C++ implementation in the proposed sgdGMF package.
@@
Appendix D Supplementary Figures and Tables
Supplementary Table 1: Exponential family laws along with their support space, canonical link ($g_{c}$), variance ($\nu$), and rescaled deviance ($D/2a$) functions. Here, we denote by $\alpha>0$ the variance parameter of the Negative Binomial distribution.
Distribution Support $g_{c}(\mu)$ $\nu(\mu)$ $D(y,\mu)/2a(\phi)$
Gaussian $\mathbb{R}$ $\mu$ 11 $(y-\mu)^{2}/2$
Gamma $\mathbb{R}_{+}$ $1/\mu$ $\mu^{2}$ $(y-\mu)/\mu-\log(y/\mu)$
Inv. Gauss. $\mathbb{R}_{+}$ $1/\mu^{2}$ $\mu^{3}$ $(y-\mu)^{2}/(2y\mu^{2})$
Poisson $\mathbb{N}_{+}$ $\log(\mu)$ $\mu$ $y\log(y/\mu)-(y-\mu)$
Supplementary Figure 1: Comparison of tSNE embeddings using aSGD with a varying number of highest variable genes (hvg) and a varying number of latent factors. No clear differences are observed between the different embeddings from using 500 hvg on, and therefore a default choice of 500 hvg is made.
Supplementary Figure 2: Comparison of mean cell-line purity using aSGD with a varying number of highest variable genes. No clear differences are observed between the different purities from using 500 hvg on, and therefore a default choice of 500 hvg is made.
Supplementary Figure 3: Comparison of aSGD using block-subsampling with NBWaVE (Agostinis et al.,, 2022), glmPCA (Townes et al.,, 2019), and COAP (Liu and Zhong,, 2024) on the Arigoni dataset (Arigoni et al.,, 2024). All methods use a matrix rank of 15, which was suggested by model selection criteria. A) tSNE embeddings show no clear differences between all methods. B) aSGD is orders of magnitude faster compared to the other methods, and has a lower out-of-sample deviance prediction error when predicting missing values. This is probably due to NBWaVE, glmPCA, and COAP requiring imputation of missing values prior to computation of the latent structure, while aSGD can deal with missing values internally. Also, RAM memory usage is similar between aSGD and COAP, which are both much more efficient compared to NBWaVE and glmPCA. C) The mean cell-line cluster purity is also similar for all methods. aSGD gives similar results on all levels when using the Poisson and Negative Binomial family on the Arigoni dataset. Therefore, aSGD is a fast alternative to NBWaVE, glmPCA, and COAP, while having similar performance.
Supplementary Figure 4: Eigenvalues of the latent directions ordered in descending values for the TENxBrainData (Lun and Morgan,, 2023). This plot, also called an elbow-plot can be used for model selection, by looking for an elbow in the curve. Here, 10 latent factors were chosen to be used in the case study.